What is the volume , in L, of 0.257 moles of neon gas at a pressure of 1.211 atm at 579 K?
1
Expert's answer
2020-04-27T01:45:45-0400

"PV=nRT"

V=(0.257*8.314*576)/1.211

V=988.62m3

V=988.62*1000=988620L